Cellophane tape 2 rolls for $1.28
Notepads 3pads for $.87
Highlighters 1 for $1.59
Ink cartridges 2 boxes for $1.98
How much would you pay for one roll of tape, two note pads, two highlighters, and two boxes of ink cartridges?

Answer by mmm4444bot(95)

Hello There:
To find the cost of one roll of tape, we divide the given price by 2 because the given price is for twice the number of rolls that we want.
$1.28/2 = $0.64
----------------
To find the cost of two note pads, we first divide the given price by 3 because we want to know how much one note pad costs.
$0.87/3 = $0.29
Each note pad is 29 cents. Since we want to note pads, we double this amount.
$0.29*2 = $0.58
----------------
One highlighter costs $1.59. We want two of them, so we double this amount.
$1.59*2 = $3.18
----------------
We want two boxes of ink catridges, and we are told that two boxes cost $1.98.
----------------
Let's total the amounts:
$0.64 + $0.58 + $3.18 + $1.98 = $6.38
We would pay $6.38 for these seven items.
~ Mark